OpenStudy (anonymous):

x^3 + 5x^2 + 3x + 15 = 0

OpenStudy (zarkon):

\[x=-5,\pm\sqrt{3}i\]

hero (hero):

x^3+5x^2+3x+15 = 0 x^2(x+5)+3(x+5) = 0 (x+5)(x^2+3) = 0 x=-5, sqrt{3}i